OpenStudy (anonymous):

using the substitution method 9x-2y=3 and 3x-6y=y not really sure how to get this going please help

OpenStudy (anonymous):

9x-2y=3 and 3x-6y=y on the right is that really supposed to equal y?

OpenStudy (anonymous):

the problems were on top of one another and it just tells me to solve it and see if it has a solution

OpenStudy (anonymous):

ok, well then we can start with the equation on the right. 3x -6y = y 3x = 7y x = (7/3)y . (7/3)y into the first equation wherever you see x.

OpenStudy (anonymous):

using the set builder notation

OpenStudy (thesecret20111):

substitute x = (7/3)y into the first equation to get: 9(7/3y)-2y=3 then solve for y.

OpenStudy (anonymous):

thank you so much for your help

OpenStudy (thesecret20111):

No problem :D